OpenStudy (anonymous):

One circle is inscribed in, and a second circle is circumscribed about, a regular polygon of 2012 sides. If the polygon’s perimeter is 2012, what is the area of the region between the two circles?

OpenStudy (anonymous):

is it: a) \[\frac{ \pi }{ 2 }\] b)\[\frac{ \pi }{ 4 }\] c)\[\frac{ \pi }{ 6 }\] d)\[\frac{ \pi }{ 1006 }\] e)\[\frac{ \pi }{ 2012 }\]

ganeshie8 (ganeshie8):

draw a rough diagram

OpenStudy (anonymous):

i tried but it looks really weird :( and it doesn't help that the polygon has 2012 sides....

ganeshie8 (ganeshie8):

lets draw a polygon wid n sides, and work it... we can plugin n=2012 in the end

OpenStudy (anonymous):

k :)

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):